Title: Systematic Lists -Three to Make 245
Question:
The sum of three numbers is 245. The ratio of the first to the second is 2 to 3, and the ratio of the second to the third is 5 to 8. What is the second Number?
My work <----
__+0+___=245 2:3=2/3
2:3 3 times 5=15 D=Digit The second Number is 75
5:8
D1 D2 D3
10+15+24=49
20+30+48=98
30+45+72=147
40+60+96=196
50+75+120=245

Please check for me




Answer by mananth(16946)

let the numbers be x ,y,z
x/y = 2/3
x= 2y/3.....................(1)
y/z = 5/8
y= 5z/8>...................(2)
plug y in (1)
x= 2(5z/8) /3
x= 10z/24
x+y+z=245
10z/24 +5z/8+z=245
multiply equation by 24
10z+15z+24z= 245*24
49z=245*24
z= 245*24/49
z=5*24
z=120
y/z= 5/8
y/120=5/8
y=120*5/8
y=15*5
y=75
x/y = 2/3
x/75=2/3
x=75*2/3
x=50
CHECK
x/y = 50/75=2/3
y/z=75/120 = 5/8
50,75,120


you are right